OpenStudy (anonymous):

rewrite in simplest radical form. show each step of the process. x^5/6 over x^1/6

Nnesha (nnesha):

when we divide same bases we should `subtract` their exponents \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]

OpenStudy (anonymous):

so 5/6-1/6= would it be 4/6

Nnesha (nnesha):

yes reduce the fraction

OpenStudy (anonymous):

so it'd be reduced to 2/3

Nnesha (nnesha):

and then rewrite that into radical form \[\huge\rm x^\frac{ m }{ n }= \sqrt[n]{x^m}\]

Nnesha (nnesha):

that's an example^^

OpenStudy (anonymous):

so it would be ^3square root x ^2. sorry if it's confusing the way I wrote it I just don't know how to add the symbol

Nnesha (nnesha):

yes that's right you should say cube root of x^2 not square root that would be \[\sqrt{ }\]

Nnesha (nnesha):

good job!

OpenStudy (anonymous):

thank u so much

Nnesha (nnesha):

mypleasure! :=)